OpenStudy (anonymous):

what is the vertex of y=x^2-11x+28

OpenStudy (anonymous):

The vertex can be found by taking the midpoint of the two zeros (4 and 7) and pluging that number into the equation. So, the midpoint is x = 5.5 y = (5.5)^2 - 11(5.5) + 28 y = 30.25 - 60.5 + 28 y = -2.25 The vertex is (5.5, -2.25)

OpenStudy (anonymous):

\[vertex=(p,q)\] \[p=\frac{-b}{2a}....q=c-ap^2\] \[y(x)=ax^2+bx+c\]\[a=1.........b=-11........c=28\]\[p=\frac{-(-11)}{2}=\frac{11}{2}....................q=28-(1)(11/2)^2=\frac{-9}{4}\]\[vertex (\frac{11}{2},\frac{-9}{4})\]

OpenStudy (anonymous):

senoins method works but there is the possibility that x intercepts will not exist.